How many numbers from 1 to 200 inclusive are equal to the cube of an integer?
(A) 1 (B) 2 (C) 3 (D) 4 (E) 5

Answer by palanisamy(496)

1^3 = 1
2^3 = 8
3^3 = 27
4^3 = 64
5^3 = 125
6^3 = 216 > 200
So There are 5 numbers from 1 to 200 inclusive are equal to the cube of an integer